Solve for w
r=(l+3w)/2

Answer by ikdeep(189)

r=(l+3w)/2
multiply both sides by 2 ,, and we get
2r = l + 3w
subtract l from both sides ,,we get
2r - l = 3w
divide both sides by 3 and we get ,,
(2r- l)/ 3 = w
hope this will help you
Please feel free to revert back for any further queries.